OpenStudy (anonymous):

Somebody please help me asap! best answer/medal Perform the indicated operation. Be sure the answer is reduced. 9-s^2/3rs divided by s^2+2rs-15/5rs. I tried to get the answer but I could not get it right. please someone!!!! HELP

OpenStudy (anonymous):

\[9-s ^{2}/ 3rs \div s ^{2}+2s-15/5rs\]

OpenStudy (anonymous):

the answe options are

OpenStudy (anonymous):

\[9-s ^{2}/3rs *5sr/s ^{2}+2rs-15=\frac{( 3-s)(3+s) }{ 3rs }*\frac{ 5rs }{( s+5)(s-3) }=-5/3(\frac{ (s+3) }{ s+5})\]

OpenStudy (anonymous):

thanks you so much!!!!